OpenStudy (anonymous):

Taylor is making a cookie cake that has a diameter of 13 in. What is the area of a 4th of the cookie cake ? A) 36.97 in. sq. B) 36.25 C) 33.18 D) 38.4

OpenStudy (apoorvk):

radius = half diameter area of a circular form = \(\pi r^2\), where 'r' is the radius. and you need to find 1/4th the area, so, you need \(\frac 1 4 \pi r^2\). can you find this out?

OpenStudy (anonymous):

132.665 divided by 1/4... how do you divide something by 1.4 on a calculator ?

OpenStudy (anonymous):

@apoorvk

OpenStudy (apoorvk):

you mean divide by '4' right?

OpenStudy (apoorvk):

type in the the no. you want to divide, and then press '/', and then 4 and then '='. or just multiply by 0.25, as 1/4 = 0.25

OpenStudy (anonymous):

so C ?

OpenStudy (apoorvk):

Yes, right-o!!

OpenStudy (apoorvk):

@Senior012